# Length-Induced Embedding Collapse in Transformer-based Models

### v1

@@ Appendix D Embedding Models
Appendix E More Experiments and Analysis
E.1 Rewriting Process
To investigate the differences in embedding distributions across texts of varying lengths, we use the Llama3 (i.e., Llama-3.1-8B-Instruct(Dubey et al., 2024) model to rewrite the texts. Specifically, we used two prompts, “Please express the given text in one sentence. No more than 10 tokens. {{original text}}” and “Please use few sentences to summarize the given text. {{original text}}”, to summarize the texts. This rewriting allows the texts to retain the same semantics while having lower lengths. By studying the differences in texts of varying lengths, we conclud that the cause of the length collapse phenomenon is that longer texts cluster to each other in embedding space.
E.2 Details on Figure 2
Figure 10: Visualization of the intensity of high-frequency components and their theoretical upper bounds. The setting is the same with Figure 2.
To verify Theorem 2, we illustrate the high-frequency intensity of each layer’s output along with its theoretical upper limit. Our visualization is based on the official checkpoint of 12-layer ANCE (Xiong et al., 2021), GIST (Solatorio, 2024) and BGE (Chen et al., 2024). We use a logarithmic scale for the purpose of a better view. Let $\bm{X}_{l}$ denote the output of the $l$-th layer. For red line, we directly calculate $\log(\lVert\mathcal{HC}\left[{\bm{X}_{l+1}}\right]\rVert_{F}/\lVert\mathcal{HC}\left[{\bm{X}_{l}}\right]\rVert_{F})$ at each layer. In practice, models typically employ multi-head attention, so we replace $\lVert\bm{W}_{V}\rVert_{2}$ in Eqn. 2 with $\sigma_{1}H$, where $\sigma_{1}=\max_{h=1}^{H}\lVert\bm{W}_{V}^{h}\rVert_{2}$. For blue line, we obtain the coefficient $\sigma_{1}$ with respect to network parameters and apply the logarithmic scale. To summarize, Figure 2 imply an convergence rate, which is consistent with our Theorem 2. Figure 10 show the same trend as Figure 2, although the E5 model exhibits anomalies at deeper layers.
E.3 Details on Figure 5
(a) ANCE
(b) GTR
(c) GSIT
(d) E5
To verify Theorem 3, we visualize the value of $\sigma_{a}$ across different text length. Our visualization is based on the texts from NFCorpus. We sample 100 samples for each bin from 0 to 500 with a bin size of 50. The $\sigma_{a}$ value is computed as the average of the $\sigma_{a}$ based on the attention of all heads before the output of the final layer. To summarize, Figure 5 imply that $sigma_{s}$ shows a decreasing trend as $n$ increases, which is consistent with our Theorem 3. Moreover, $\sigma_{a}$ also increases as $\tau$ decreases, further validating our proposed method, TempScale. Figure 11 also show the same trend as Figure 5 and the E5 model shows an increasing $\sigma_{a}$ with length, which may be due to anomalies in the deep layer attention patterns.
E.4 More Analysis about Assumption in Theorem 4
Figure 12: The cosine similarity of embeddings of texts generated by repeating word “dog” and “cat”.
In Corollary 4, we hypothesize and verify that all natural language sequences tend to have a relatively consistent representation. As a result, different texts tend to exhibit consistent low-pass signals after losing high-frequency information, leading to ultimately consistent text embeddings. This leads to an increase in cosine similarity for longer texts. However, as shown in Figure 12, we repeat the word “dog” and “word” $n$ times and calculate the similarity between these two texts. The results show that even when the sequences do not have overlap, the text embeddings tend to converge to similar representations as the sequence length increases. This further demonstrates that length collapse causes completely different sequences to converge toward similarity.
E.5 More Results on Longer Queries
Figure 13: More results about performance difference between long query and short query across varying temperature $\tau$ on SummScreenFD dataset.
In previous experiments, we find that lower temperatures improve the performance of long queries on the SummScreenFD datset. In Figure 13, we observe a similar phenomenon across other models. Except for ANCE, the performance of other models on long queries decreases as the temperature decreases. This suggests that long queries require a lower temperature to mitigate the length collapse phenomenon. However, ANCE’s performance degradation with decreasing temperature is likely attributable to its inherent limitations in processing long texts.
E.6 More Results about Temperature Changes with Length
Figure 14: More results about performance difference between long query and short query across varying temperature $\tau$ on AmazonPolarity dataset.
Figure 15: Relative performance compared to the raw results with varying query (Q Temperature) and document (D Temperature) temperatures using different model on SummScreenFD
As shown in Figure 14, texts of different length intervals perform best at different temperatures $\tau$. In general, short texts perform better at higher $\tau$, while longer texts require a lower $\tau$ to prevent length collapse. This suggests we can set different $\tau$ values for texts in different length intervals to achieve better overall performance. Moreover, as shown in Figure 15, compared to the QMSum dataset, SummScreenFD requires a lower temperature for scaling the document due to its longer length. This further supports the conclusion that longer texts require a lower temperature for scaling.

### v2

@@ Appendix F Datasets and Evaluation Metrics
Table 9: Embedding models used in the experiments.
F.1 Classification
AmazonPolarity Zhang et al. (2015) consists of Amazon customer reviews, each labeled as either “positive” or “negative.”
Banking77  Casanueva et al. (2020) dataset consists of online banking user queries labeled with one of 77 specific intents.
Emotion  Saravia et al. (2018) comprises Twitter messages categorized by six fundamental emotions: anger, fear, joy, love, sadness, and surprise.
IMDb  Maas et al. (2011) consists of extensive movie reviews categorized as either positive or negative.
MassiveIntent  FitzGerald et al. (2022) is a multilingual dataset featuring a diverse array of utterances from Amazon Alexa, each labeled with one of 60 different intents across 51 languages.
MassiveScenario  FitzGerald et al. (2022) dataset comprises a diverse collection of Amazon Alexa user utterances, each labeled with one of 60 thematic intents, and supports 51 languages.
TweetSentimentExtraction  [^2] TweetSentimentExtraction, a dataset from a Kaggle competition, focuses on classifying tweets into three categories: neutral, positive, and negative sentiments.
F.2 Clustering
Clustering aims at grouping a given set of sentences or texts into meaningful clusters by training a mini-batch k-means model on the text embeddings. The model is scored using the v-measure  Rosenberg and Hirschberg (2007). Since the V-measure does not depend on the cluster labels, the arrangement of the labels will not affect the score.
ArxivClusteringS2S, ArxivClusteringP2P, BiorxivClusteringS2S, BiorxivClusteringP2P, MedrxivClusteringP2P, MedrxivClusteringS2S Muennighoff et al. (2023). These datasets are tailored for MTEB, utilizing titles or a combination of titles and abstracts from arXiv, bioRxiv, and medRxiv, with clustering labels derived from human-assigned categories, emphasizing both main and secondary classification levels.
RedditClustering Geigle et al. (2021), a dataset that consists of titles from 199 subreddits, is organized into 25 splits, each featuring 10 to 50 classes, with every class containing between 100 and 1,000 sentences.
RedditClusteringP2P Muennighoff et al. (2023), developed for the MTEB, consists of Reddit posts combined with their titles, organized into ten splits featuring 10 and 100 clusters each, with a total of 1,000 to 100,000 posts, aimed at clustering based on subreddit affiliation.
StackExchangeClustering Geigle et al. (2021), a dataset consisting of titles from 121 Stack Exchange communities, is organized into 25 subsets, each containing 10 to 50 categories, with 100 to 1,000 sentences per category.
StackExchangeClusteringP2P Muennighoff et al. (2023), designed for MTEB, comprises 10 splits of posts from StackExchange, each containing 5,000 to 10,000 entries, clustered by subreddit based on the combined content of titles and posts.
F.3 Reranking
MindSmall Wu et al. (2020) dataset is a comprehensive English resource designed for research in news recommendation, focusing on ranking news articles based on the title of a currently read article to suggest related content.
SciDocsRR Cohan et al. (2020a) is a dataset designed for ranking related scientific papers using their titles as the primary basis for assessment.
StackOverflowDupQuestions Liu et al. (2018) dataset focuses on identifying whether questions tagged with Java, JavaScript, and Python on Stack Overflow are duplicates of existing queries.
F.4 Retrieval
In the retrieval task, each dataset consists of a corpus, queries, and a mapping of each query to relevant documents. The task goal is to find these relevant documents based on a given query. When evaluating, we first use the provided model to embed queries and corpus documents and then calculate the cosine similarity to obtain relevance scores and rank the corpus documents for each query based on these scores. The evaluation metrics consist of nDCG@k, MRR@k, MAP@k, precision@k, and recall@k, with nDCG@10 as the primary metric.
F.4.1 Beir Retrival
NFCorpus Boteva et al. (2016) is a dataset that includes natural language queries sourced from NutritionFacts, paired with annotated medical documents from PubMed, utilizing the original splits from various types of content from NF, such as videos, blogs, and Q&A posts.
SciFact Wadden et al. (2020) dataset evaluates scientific claims by matching them with evidence sourced from research literature, specifically utilizing a set of 300 test queries and the complete document collection from the original dataset.
SCIDOCS Cohan et al. (2020b) is a benchmark dataset for evaluating scientific document embeddings, featuring seven tasks such as citation prediction, document classification, and recommendation.
FiQA2018 Maia et al. (2018) is a financial question answering dataset built by crawling StackExchange posts under the Investment topic from 2009 to 2017, containing a knowledge base of 57,640 answer posts and 17,110 training question-answer pairs, with 531 testing question-answer pairs.
Touche2020 Bondarenko et al. (2020) is a benchmark for argument retrieval, designed to support research in finding and evaluating arguments on various topics.
F.4.2 LongEmbd Retrival
LongEmbed Zhu et al. (2024) includes 4 real-world retrieval tasks curated from long-form QA and summarization. The document in LongEmbed is much longer compared to BEIR. Thus, it can effectively evaluate the capability of the embedding model on long texts.
LEMBNarrativeQARetrieval Kočiskỳ et al. (2018) is a question-answering dataset featuring lengthy narratives, averaging 50,474 words, that challenge models to comprehend and extract information about characters and events dispersed throughout the stories.
LEMBQMSumRetrieval Zhong et al. (2021) dataset focuses on generating summaries of meetings based on specific queries, necessitating the extraction and synthesis of relevant information from various segments of the conversation that cover multiple topics and participants.
LEMBSummScreenFDRetrieval Chen et al. (2021) dataset consists of pairs of transcripts from TV series and their corresponding human-crafted summaries, requiring the integration of dispersed plot elements into concise narrative descriptions.
LEMBWikimQARetrieval Ho et al. (2020) dataset is a complex question-answering resource that includes questions requiring up to five reasoning steps, designed using specific templates to encourage deep understanding rather than simple retrieval of information.
F.5 Semantic Textual Similarity (STS)
The task goal is to determine the similarity between a pair of sentences, where continuous scores serve as labels, with higher values indicating greater similarity. The provided model embeds the sentences, and their similarity is calculated using cosine similarity. The primary evaluation metric is the Spearman correlation Reimers et al. (2016).
STS12, STS13, STS14, STS15, STS16, STS17, STS22, STSBenchmark Agirre et al. (2012, 2013); Bandhakavi et al. (2014); Biçici (2015); Nakov et al. (2016)  [^5] https://alt.qcri.org/semeval2017/task1/  [^6] https://competitions.codalab.org/competitions/33835  [^7] https://github.com/PhilipMay/stsb-multi-mt/ are collections of sentence pairs designed to evaluate semantic textual similarity, with the former set focused on monolingual English pairs and the latter two incorporating cross-lingual comparisons across multiple languages.
BIOSSES Soğancıoğlu et al. (2017) comprises 100 pairs of sentences specifically focused on the biomedical domain.
SICK-R Dadas et al. (2020), which stands for Sentences Involving Compositional Knowledge, comprises 100,000 diverse sentence pairs that exhibit rich lexical, syntactic, and semantic characteristics.
F.6 Summarization
The input consists of a set of summaries written by humans and machines. The goal is to score the machine-generated summaries. Use the provided model to embed the summaries. Calculate the distance between each machine summary and all human summary embeddings. Retain similar scores to the model score for each individual machine-generated summary. Calculate the Spearman correlation based on cosine similarity Reimers et al. (2016) as the main metric.
SummEval Fabbri et al. (2021) consists of summaries produced by advanced summarization models trained on CNN and DailyMail articles.
Appendix G Embedding Models
ANCE Xiong et al. (2021) enhances dense retrieval by selecting challenging negative samples from the entire corpus and asynchronously updating the Approximate Nearest Neighbor (ANN) index with each training iteration, using a context window size of 512.
GTR Ni et al. (2022) improves dual encoder performance for retrieval tasks by scaling up model size while keeping a fixed bottleneck embedding, leading to significant improvements in out-of-domain generalization, all within a context window size of 512.
GIST Solatorio (2024) consistently improves performance across different model sizes by leveraging the strengths of large, resource-intensive models to enhance smaller ones, making advanced AI technologies more accessible and cost-effective, all within a context window size of 512.
BGE Xiao et al. (2023) offers a range of well-trained embedding models based on a BERT-like architecture, enabling users to balance performance and efficiency for various applications while also allowing easy fine-tuning. In our experiments, we use the gte-base-en-v1.5 model, which operates with a context window size of 512.
E5 Zhu et al. (2024) is a long-context embedding model fine-tuned to support 4k token inputs while maintaining the original performance for shorter contexts, designed to advance research in long-context embedding technologies. It uses a context window size of 4k.